OpenStudy (anonymous):

find the energy in KWh consumed in 10hrs by four devices of power 500 w each.

OpenStudy (anonymous):

0.5KW * 10 = 5KWH*4 =20 KWH

OpenStudy (anonymous):

t = 10 hrs Power of 1 device = 500 watts Therefore,Power of 4 devices = 4*500 = 2000 watts 1000 watts = 1 Kw Therefore, 2000 watts= 2000/1000 = 2Kw Power=Energy/Time , which implies , Energy = Power x Time Therefore, Energy (in Kwh) = 2 x 10 =20 Kwh